# A Framework for Safe Probabilistic Invariance Verification of Stochastic Dynamical Systems

### v1

@@ 4 Safe Probabilistic Invariance Verification for Discrete-time Systems
This section presents two sets of optimizations to addressing the safe probabilistic invariance verification problem in Definition 2. The first set of optimizations is adapted from stochastic barrier certificates for safety and reachability verification in [3]. The second set of optimizations is inspired by Theorem 1 along with Proposition 1.
@@
$\begin{cases}\bm{x}(j+1)=\widehat{\bm{f}}(\bm{x}(j),\bm{d}(j))&\forall j\in\mathbb{N}\\ \bm{x}(0)=\bm{x}_{0}\end{cases},$ (5)
@@
$\widehat{\mathcal{X}}\supset\{\bm{x}\in\mathbb{R}^{n}\mid\bm{x}=\bm{f}(\bm{x}_{0},\bm{d}),\bm{x}_{0}\in\mathcal{X},\bm{d}\in\mathcal{D}\}\cup\mathcal{X}.$ (6)
@@
Also, since $\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(1)=\bm{\phi}_{\pi}^{\bm{x}_{0}}(1)$ for $\bm{x}\in\mathcal{X}$, we have that
@@
Given a disturbance signal $\pi$ and an initial state $\bm{x}_{0}\in\mathcal{X}$, the resulting trajectory $\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(\cdot)\colon\mathbb{N}\rightarrow\mathbb{R}^{n}$ either enters the unsafe set $\widehat{\mathcal{X}}\setminus\mathcal{X}$ in finite time (i.e., $\exists k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\wedge\forall i\in\mathbb{N}_{\leq k-1}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(i)\in\mathcal{X}$ ) or stays inside the safe set $\mathcal{X}$ always (i.e., $\forall k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}$). Thus,
$\begin{split}&\mathbb{P}^{\infty}(\exists k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\wedge\forall i\in\mathbb{N}_{\leq k-1}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(i)\in\mathcal{X})\\ &+\mathbb{P}^{\infty}(\forall k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X})=1.\end{split}$
Therefore, if the upper and lower bounds of the probability $\mathbb{P}^{\infty}(\exists k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\wedge\forall i\in\mathbb{N}_{\leq k-1}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(i)\in\mathcal{X})$ are gained, one can obtain the lower and upper bounds of the liveness probability of staying inside the safe set $\mathcal{X}$.
4.1 Barrier Certificates Based Invariance Verification
In this subsection we propose optimizations for certifying lower and upper bounds of the liveness probability by adapting barrier certificates in [3] for safety and reachability verification, which are respectively built upon the well established Doob’s nonnegative supermartingale inequality [26].
Proposition 2 provides a straightforward sufficient condition for lower bounds on the liveness probability, as demonstrated in Theorem 5 of [3].
Proposition 2
@@ 4.1 Barrier Certificates Based Invariance Verification
$\begin{cases}v(\bm{x})\leq 1-\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\mathcal{X}_{unsafe}(=\Omega\setminus\mathcal{X})\\ \mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}})]-v(\bm{x})\leq 0&\forall\bm{x}\in\Omega\end{cases},$ (7)
then $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{unsafe}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\leq 1-\epsilon_{1}$. Thus, $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\geq\epsilon_{1}$.
Similarly, a sufficient condition for determining upper bounds of the liveness probability can be obtained straightforwardly from Theorem 16 in [3]. However, finding a robust invariant set $\Omega$, except for the trivial case of $\Omega=\mathbb{R}^{n}$, can be challenging and computationally intensive for many systems, if it even exists. However, when $\Omega=\mathbb{R}^{n}$ in Proposition 2, the resulting constraint (7) may be too strong, leading to an overly conservative lower bound. We use an example to illustrate this below.
Example 1
@@
$\begin{cases}x(l+1)=x(l)-0.02y(l)\\ y(l+1)=y(l)+0.01\Big{(}(0.8+d(l))x(l)\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+10(x^{2}(l)-0.21)y(l)\Big{)}\end{cases},$ (8)
where $d(\cdot)\colon\mathbb{N}\rightarrow\mathcal{D}=[-0.1,0.1]$, $\mathcal{X}=\{\,(x,y)^{\top}\mid h(\bm{x})\leq 0\,\}$ with $h(\bm{x})=x^{2}+y^{2}-1$, and $\mathcal{X}_{0}=\{\,(x,y)^{\top}\mid g(\bm{x})<0\,\}$ with $g(\bm{x})=x^{2}+y^{2}-0.01$.
Via solving Op0, which is encoded into a semi-definite program SDP0 (shown in Appendix) via the sum of squares decomposition for multivariate polynomials:
$\begin{split}&\textbf{Op0}~{}~{}~{}~{}~{}~{}~{}\max_{v(\bm{x}),\epsilon_{1}}\epsilon_{1}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})\leq 1-\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ v(\bm{x})\geq\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]&\forall\bm{x}\in\mathbb{R}^{n}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\mathbb{R}^{n}\setminus\mathcal{X}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\mathbb{R}^{n},\\ \epsilon_{1}\geq 0\end{cases}.\end{split}$
we obtain a lower bound of the liveness probability, which is 2.1368e-07. This is too conservative to be useful in practice. The resulting semi-definite program is addressed when unknown polynomials of degree 8 are used. $\blacksquare$
In the following we will present weaker sufficient conditions for certifying lower and upper bounds using the switched system (5). They are respectively formulated in Proposition 3 and 4.4.
Proposition 3
@@
$\begin{cases}v(\bm{x})\leq 1-\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ v(\bm{x})\geq\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\widehat{\mathcal{X}}\end{cases},$ (9)
then $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\leq 1-\epsilon_{1}$. Consequently, $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\geq\epsilon_{1}$.
Proof 4.2.
@@
Therefore, $v(\bm{x})$ is a classical nonnegative supermartingale based barrier certificate for system (5) with the invariant set $\widehat{\mathcal{X}}$ and the unsafe set $\widehat{\mathcal{X}}\setminus\mathcal{X}$. Therefore, according to Theorem 5 in [3], we have the conclusion that the probability of reaching the unsafe set $\widehat{\mathcal{X}}\setminus\mathcal{X}$ for system (5) starting from $\mathcal{X}_{0}$ is less than or equal to $1-\epsilon_{1}$, i.e., $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\leq 1-\epsilon_{1}$. Therefore,
$\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\geq\epsilon_{1}.$
Since if $\bm{x}_{0}\in\mathcal{X}$, $\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(1)=\bm{\phi}^{\bm{x}_{0}}_{\pi}(1)$ holds. Consequently, $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\geq\epsilon_{1}$.
According to Proposition 3, a lower bound of the liveness probability can be computed via solving the following optimization:
$\begin{split}&\textbf{Op1}~{}~{}~{}~{}~{}~{}~{}\max_{v(\bm{x}),\epsilon_{1}}\epsilon_{1}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})\leq 1-\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ v(\bm{x})\geq\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\widehat{\mathcal{X}}\\ \epsilon_{1}\geq 0\end{cases}.\end{split}$
@@
Consider Example 1 once again. By solving Op1 using $\widehat{\mathcal{X}}=\{\,(x,y)^{\top}\mid\widehat{h}(\bm{x})\leq 0\,\}$ with $\widehat{h}(\bm{x})=x^{2}+y^{2}-2$, we encode it into a semi-definite program SDP1(shown in Appendix) via the sum of squares decomposition for multivariate polynomials. The solution yields a lower bound for the liveness probability, which is 0.9465. SDP1 is addressed when unknown polynomials of degree 8 are used. $\blacksquare$
Proposition 4.4.
@@
$\begin{cases}v(\bm{x})\leq\epsilon_{2}&\forall\bm{x}\in\mathcal{X}_{0},\\ v(\bm{x})\geq 1&\forall\bm{x}\in\partial\widehat{\mathcal{X}}\setminus\partial\mathcal{X}_{unsafe}\\ \mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]-v(\bm{x})\leq-\delta&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\widehat{\mathcal{X}}\end{cases},$ (10)
where $\delta>0$ is a user-defined value, then $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\leq\epsilon_{2}$.
Proof 4.5.
@@
$\begin{cases}v(\bm{x})\leq\epsilon_{2}&\forall\bm{x}\in\mathcal{X}_{0},\\ v(\bm{x})\geq 1&\forall\bm{x}\in\partial\widehat{\mathcal{X}}\setminus\partial\mathcal{X}_{unsafe}\\ \mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}}(1))]-v(\bm{x})\leq-\delta&\forall\bm{x}\in\overline{\widehat{\mathcal{X}}\setminus\mathcal{X}_{unsafe}}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\widehat{\mathcal{X}}\end{cases},$ (11)
According to Theorem 16 in [3] and following the proof of Proposition 3, we have the conclusion $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\leq\epsilon_{2}$.
@@
If $v(\bm{x})$ is bounded over $\widehat{\mathcal{X}}$ in (10), constraint (10) provides strong guarantees of leaving the safe set $\mathcal{X}$ almost surely, i.e., $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}=0$. This conclusion is justified as follows.
@@
Thus, for $\bm{x}\in\widehat{\mathcal{X}}$, we have that
@@
which implies $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}\in\mathcal{X}\big{)}=\lim_{k\rightarrow\infty}\frac{\sum_{i=0}^{k-1}\mathbb{E}^{\infty}[1_{\mathcal{X}_{unsafe}}(\widehat{\bm{\phi}}_{\pi}^{\bm{x}}(i))]}{k}\geq 1$ (according to Lemma 3 in [29]) Thus, we have the conclusion.
It is worth noting that if $\partial\widehat{\mathcal{X}}\cap\partial\mathcal{X}=\emptyset$, the set $\partial\widehat{\mathcal{X}}\setminus\partial\mathcal{X}_{unsafe}$ in (10) is empty. As a result, the constraint $v(\bm{x})\geq 1,\forall\bm{x}\in\partial\widehat{\mathcal{X}}\setminus\partial\mathcal{X}_{unsafe}$ becomes redundant and can be removed. Throughout this paper, unless explicitly stated otherwise, we assume that $\partial\widehat{\mathcal{X}}\cap\partial\mathcal{X}=\emptyset$. It is worth mentioning that this assumption is not overly strict and can be easily satisfied by enlarging the set that satisfies (6). The primary role of this assumption is to facilitate solving constraint (10). Accordingly, based on Proposition 4.4, we can calculate an upper bound for the liveness probability by solving the following optimization problem:
$\begin{split}&\textbf{Op2}~{}~{}~{}~{}~{}~{}~{}\min_{v(\bm{x}),\epsilon_{2}}\epsilon_{2}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})\leq\epsilon_{2}&\forall\bm{x}\in\mathcal{X}_{0}\\ \mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]-v(\bm{x})\leq-\delta&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\widehat{\mathcal{X}}\\ \epsilon_{2}\geq 0\end{cases}.\end{split}$
@@
Another condition, which is analogous to the one in Proposition 4.4, was proposed in [5].
@@
$\begin{cases}v(\bm{x})\geq c&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})<c&\forall\bm{x}\in\mathcal{X}_{unsafe}(=\widehat{\mathcal{X}}\setminus\mathcal{X})\\ \mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]-v(\bm{x})\leq-1&\forall\bm{x}\in\mathcal{X}\end{cases},$ (12)
then $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}\in\mathcal{X}_{0}\big{)}=0$.
Proof 4.9.
The proof is similar to the one of Proposition 3, and the conclusion is justified from Theorem 19 in [3].

### v2

@@ 4 Safe Probabilistic Invariance Verification for Discrete-time Systems
This section presents two sets of optimizations to addressing the safe probabilistic invariance verification problem in Definition 3.2. The first set of optimizations is adapted from stochastic barrier certificates for safety and reachability verification in [3]. The second set of optimizations is inspired by Proposition 3.4 along with Corollary 3.5.
@@
(5) $\begin{cases}\bm{x}(j+1)=\widehat{\bm{f}}(\bm{x}(j),\bm{d}(j))&\forall j\in\mathbb{N}\\ \bm{x}(0)=\bm{x}_{0}\end{cases},$
@@
(6) $\widehat{\mathcal{X}}\supset\{\bm{x}\in\mathbb{R}^{n}\mid\bm{x}=\bm{f}(\bm{x}_{0},\bm{d}),\bm{x}_{0}\in\mathcal{X},\bm{d}\in\mathcal{D}\}\cup\mathcal{X}.$
@@
Also, since $\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(1)=\bm{\phi}_{\pi}^{\bm{x}_{0}}(1)$ for $\bm{x}_{0}\in\mathcal{X}$, we have that
@@
Given a disturbance signal $\pi$ and an initial state $\bm{x}_{0}\in\mathcal{X}$, the resulting trajectory $\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(\cdot)\colon\mathbb{N}\rightarrow\mathbb{R}^{n}$ either enters the unsafe set $\widehat{\mathcal{X}}\setminus\mathcal{X}$ in finite time (i.e., $\exists k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\wedge\forall i\in\mathbb{N}_{\leq k-1}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(i)\in\mathcal{X}$ ) or stays inside the safe set $\mathcal{X}$ always (i.e., $\forall k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}$). Thus, $\mathbb{P}^{\infty}(\exists k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\wedge\forall i\in\mathbb{N}_{\leq k-1}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(i)\in\mathcal{X})+\mathbb{P}^{\infty}(\forall k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X})=1$. Therefore, if the upper and lower bounds of the probability $\mathbb{P}^{\infty}(\exists k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\wedge\forall i\in\mathbb{N}_{\leq k-1}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(i)\in\mathcal{X})$ are gained, one can obtain the lower and upper bounds of the liveness probability of staying inside the safe set $\mathcal{X}$.
4.1 Doob’s Nonnegative Supermartingale Inequality Based Invariance Verification
In this subsection we propose optimizations for certifying lower and upper bounds of the liveness probabilities by adapting barrier certificates in [3] for safety and reachability verification, which are respectively built upon the well established Doob’s nonnegative supermartingale inequality [27].
Proposition 4.1 provides a straightforward sufficient condition for lower bounds on the liveness probabilitiess, as demonstrated in Theorem 5 of [3].
Proposition 4.1.
@@ 4.1 Doob’s Nonnegative Supermartingale Inequality Based Invariance Verification
(7) $\begin{cases}v(\bm{x})\leq 1-\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\mathcal{X}_{unsafe}(=\Omega\setminus\mathcal{X})\\ \mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]-v(\bm{x})\leq 0&\forall\bm{x}\in\Omega\end{cases},$
then $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{unsafe}\big{)}\leq 1-\epsilon_{1}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$. Thus, $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\big{)}\geq\epsilon_{1}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$
Similarly, a sufficient condition for determining upper bounds of the liveness probabilities can be obtained straightforwardly from Theorem 16 in [3]. However, finding a robust invariant set $\Omega$, except for the trivial case of $\Omega=\mathbb{R}^{n}$, can be challenging and computationally intensive for many systems, if it even exists. On the other hand, when $\Omega=\mathbb{R}^{n}$ in Proposition 4.1, the resulting constraint (7) may be too strong, leading to an overly conservative lower bound. We use an example to illustrate this below.
Example 4.2.
@@
(8) $\begin{cases}x(l+1)=x(l)+0.01\Big{(}-2y(l)\Big{)}\\ y(l+1)=y(l)+0.01\Big{(}(0.8+d(l))x(l)+10(x^{2}(l)-0.21)y(l)\Big{)}\end{cases},$
where $d(\cdot)\colon\mathbb{N}\rightarrow\mathcal{D}=[-0.1,0.1]$, $\mathcal{X}=\{\,(x,y)^{\top}\mid h(\bm{x})\leq 0\,\}$ with $h(\bm{x})=x^{2}+y^{2}-1$, and $\mathcal{X}_{0}=\{\,(x,y)^{\top}\mid g(\bm{x})<0\,\}$ with $g(\bm{x})=x^{2}+y^{2}-0.01$. We assume that the probability distribution on $\mathcal{D}$ is the uniform distribution. The lower bound of the liveness probabilities estimated via the Monte Carlo method is 1.
Via solving Op0 (shown in Appendix B), which is encoded into a semi-definite program SDP0 (shown in Appendix C) via the sum of squares decomposition for multivariate polynomials, we obtain a lower bound of the liveness probabilities, which is 2.1368e-07. This is too conservative to be useful in practice. The resulting semi-definite program is addressed when unknown polynomials of degree 8 are used. $\blacksquare$
In the following we will present weaker sufficient conditions for certifying lower and upper bounds using the switched system (5). They are respectively formulated in Theorem 4.3 and 4.6.
Theorem 4.3.
@@
(9) $\begin{cases}v(\bm{x})\leq 1-\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ v(\bm{x})\geq\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\widehat{\mathcal{X}}\end{cases},$
then $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\big{)}\leq 1-\epsilon_{1}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$. Consequently, $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\big{)}\geq\epsilon_{1}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$.
Proof 4.4.
@@
Therefore, $v(\bm{x})$ is a classical nonnegative supermartingale based barrier certificate for system (5) with the invariant set $\widehat{\mathcal{X}}$ and the unsafe set $\widehat{\mathcal{X}}\setminus\mathcal{X}$. Therefore, according to Theorem 5 in [3], we have the conclusion that the probability of reaching the unsafe set $\widehat{\mathcal{X}}\setminus\mathcal{X}$ for system (5) starting from each state in $\mathcal{X}_{0}$ is less than or equal to $1-\epsilon_{1}$, i.e., $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\big{)}\leq 1-\epsilon_{1}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$. Therefore,
$\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\big{)}\geq\epsilon_{1},\forall\bm{x}_{0}\in\mathcal{X}_{0}.$
Since if $\bm{x}_{0}\in\mathcal{X}$, $\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(1)=\bm{\phi}^{\bm{x}_{0}}_{\pi}(1)$ holds. Consequently, $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\big{)}\geq\epsilon_{1}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$.
According to Theorem 4.3, a lower bound of the liveness probabilities can be computed via solving Op1 (shown in Appendix B).
@@
Consider Example 4.2 again. By solving Op1 using $\widehat{\mathcal{X}}=\{\,(x,y)^{\top}\mid\widehat{h}(\bm{x})\leq 0\,\}$ with $\widehat{h}(\bm{x})=x^{2}+y^{2}-2$, we encode it into a semi-definite program SDP1(shown in Appendix) via the sum of squares decomposition for multivariate polynomials. The solution yields a lower bound for the liveness probabilities, which is 0.9465. SDP1 is addressed when unknown polynomials of degree 8 are used. $\blacksquare$
Theorem 4.6.
@@
(10) $\begin{cases}v(\bm{x})\leq\epsilon_{2}&\forall\bm{x}\in\mathcal{X}_{0},\\ v(\bm{x})\geq 1&\forall\bm{x}\in\partial\widehat{\mathcal{X}}\setminus\partial\mathcal{X}_{unsafe}\\ \mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]-v(\bm{x})\leq-\delta&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\widehat{\mathcal{X}}\end{cases},$
where $\delta>0$ is a user-defined value, then $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\big{)}\leq\epsilon_{2}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$.
Proof 4.7.
@@
(11) $\begin{cases}v(\bm{x})\leq\epsilon_{2}&\forall\bm{x}\in\mathcal{X}_{0},\\ v(\bm{x})\geq 1&\forall\bm{x}\in\partial\widehat{\mathcal{X}}\setminus\partial\mathcal{X}_{unsafe}\\ \mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}}(1))]-v(\bm{x})\leq-\delta&\forall\bm{x}\in\overline{\widehat{\mathcal{X}}\setminus\mathcal{X}_{unsafe}}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\widehat{\mathcal{X}}\end{cases},$
According to Theorem 16 in [3] and following the proof of Theorem 4.3, we have the conclusion $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\big{)}\leq\epsilon_{2}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$.
@@
If $v(\bm{x})$ is bounded over $\widehat{\mathcal{X}}$ in (10), constraint (10) provides strong guarantees of leaving the safe set $\mathcal{X}$ almost surely, i.e., $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\big{)}=0$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$. This conclusion is justified as follows.
@@
Thus, for $\bm{x}\in\widehat{\mathcal{X}}$, we have
@@
which implies $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\big{)}=\lim_{k\rightarrow\infty}\frac{\sum_{i=0}^{k-1}\mathbb{E}^{\infty}[1_{\mathcal{X}_{unsafe}}(\widehat{\bm{\phi}}_{\pi}^{\bm{x}}(i))]}{k}$ $\geq 1$ (according to Lemma 3 in [30]) Thus, we have the conclusion.
It is worth noting that if $\partial\widehat{\mathcal{X}}\cap\partial\mathcal{X}=\emptyset$, the set $\partial\widehat{\mathcal{X}}\setminus\partial\mathcal{X}_{unsafe}$ in (10) is empty. As a result, the constraint $v(\bm{x})\geq 1,\forall\bm{x}\in\partial\widehat{\mathcal{X}}\setminus\partial\mathcal{X}_{unsafe}$ becomes redundant and can be removed. Throughout this paper, unless explicitly stated otherwise, we assume that $\partial\widehat{\mathcal{X}}\cap\partial\mathcal{X}=\emptyset$. It is worth mentioning that this assumption is not overly strict and can be easily satisfied by enlarging the set that satisfies (6). The primary role of this assumption is to facilitate solving constraint (10). Accordingly, based on Theorem 4.6, we can calculate an upper bound for the liveness probabilities by solving Op2 (shown in Appendix B).
@@
Another condition, which is analogous to the one in Theorem 4.6, was proposed in [5].
@@
(12) $\begin{cases}v(\bm{x})\geq c&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})<c&\forall\bm{x}\in\mathcal{X}_{unsafe}(=\widehat{\mathcal{X}}\setminus\mathcal{X})\\ \mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]-v(\bm{x})\leq-1&\forall\bm{x}\in\mathcal{X}\end{cases},$
then $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\big{)}=0$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$.
Proof 4.11.
The proof is similar to the one of Theorem 4.3, and the conclusion is justified from Theorem 19 in [3].